OpenStudy (anonymous):

A rectangle has vertices whose coordinates are A(1, 4), B(4, -1), C(-1, -4), and D(-4, 1). Find the coordinates of the midpoints of each side. (Enter answers as fractions.) Midpoint of AB = ( , ) Midpoint of BC = ( , ) Midpoint of CD = ( , ) Midpoint of AD = ( , )

OpenStudy (amistre64):

add AB and divide it in half to fine the average, or middle, point

OpenStudy (amistre64):

same for the rest of them

OpenStudy (amistre64):

for example: B = 4 -1 +C=-1 -4 ---------- 3 -5 mid point of BC is then 3/2 -5/2